OpenStudy (anonymous):

solve the system using the equivalent system: -2x-y=0 and 3x-y=10 the options are (4,-2), (-2,4), (2,-4), (-4,2)

OpenStudy (anonymous):

help please!

OpenStudy (anonymous):

(2,-4) -2*(2)-(-4)=-4+4=0 3*(2)-(-4)= 6+4=10

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

it easy question you should to plug the choices but if you dont have any choices you should to solve this system

OpenStudy (anonymous):

oh, so next time just plug them in?

OpenStudy (anonymous):

yes